OpenStudy (anonymous):

@ is an operation that says a@b=a(b-a),if x-y=1,x+y=2,calculate (y@x)@y

OpenStudy (anonymous):

then solve for x and y first x=? y=?

hartnn (hartnn):

0

OpenStudy (anonymous):

\[y(x-y)@y\] \[y@y=(y-y)y\]=0

hartnn (hartnn):

ya,did mentally, as u get y@y, u don't need to do anything else..